# matrix,arrows,decorations.pathmorphing

### v2

@@ 4.1. Preliminary lemmas
The homomorphisms:
@@
    $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ and
@@
    $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$,
induced by restriction of automorphisms, are injective.
@@
(i): The statement is trivial for $n(S)\leq 1$ so that we can assume $n(S)>1$. For $S=S_{1,2}$, we have that ${\check{\Gamma}}^{\pm}(S)=\mathrm{P}{\check{\Gamma}}^{\pm}(S)\times\langle{\upsilon}\rangle$, where ${\upsilon}$ is the hyperelliptic involution of ${\Gamma}(S)$. This identity implies the stronger statement that $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$.
Thus, we can assume that the center $Z({\check{\Gamma}}^{\pm}(S))$ of ${\check{\Gamma}}^{\pm}(S)$ is trivial. Since the restriction of the homomorphism $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ to $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ is injective, the conclusion then follows from Lemma 3.3 in [6].
(ii): Wells’ exact sequence (cf. Theorem in [25]), applied to the short exact sequence $1\to\mathrm{P}{\check{\Gamma}}(S)\to\mathrm{P}{\check{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2\to 1$, implies that the kernel of the given homomorphism is contained in the subgroup of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))_{\mathrm{P}{\check{\Gamma}}(S)}$ determined by the group of homomorphisms $\operatorname{Hom}(\mathbb{Z}/2,Z(\mathrm{P}{\check{\Gamma}}(S)))$. From the explicit description of this subgroup, it follows that, even in the cases when $Z(\mathrm{P}{\check{\Gamma}}(S))$ is not trivial, the intersection of the image of $\operatorname{Hom}(\mathbb{Z}/2,Z(\mathrm{P}{\check{\Gamma}}(S)))$ in $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))_{\mathrm{P}{\check{\Gamma}}(S)}$ with $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is trivial, which implies item (ii) of the lemma. ∎
@@
There is a chain of natural inclusions:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)).$
@@
By Lemma 4.5, in order to prove Theorem 1.1, it is enough to show that the images of $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ coincide. The idea is to use Theorem 2.15. We need to consider separately three different cases.
4.3. Proof of Theorem 1.1 for $g(S)=0$
@@
By Lemma 4.2, after composing with an inner automorphism of ${\widehat{\Gamma}}(S)^{\pm}$, we can assume that a given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}(S)^{\pm})$ preserves a fixed antiholomorphic involution $\iota\in\mathrm{P}\Gamma(S)^{\pm}$. The latter has for fixed point set in the surface $S$ a separating simple closed curve containing all punctures of $S$.
@@
By Lemma 4.3, the fixed point set $C(S)^{\iota}_{0}$, for the action of $\iota$ on $C(S)_{0}$, is finite and then, by Proposition 3.6, identifies with the fixed point set ${\check{C}}(S)^{\iota}_{0}$, for the action of $\iota$ on ${\check{C}}(S)_{0}$.
@@
Let $\{\alpha,\beta\}$ be a pair of $\iota$-invariant simple closed curves on $S$ which intersect precisely in two points. An element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$, such that $f(\iota)=\iota$, then preserves $C(S)^{\iota}_{0}={\check{C}}(S)^{\iota}_{0}$ and sends the pair $\{[\alpha],[\beta]\}$ of 00-simplices in $C(S)^{\iota}_{0}$ to the pair $\{f([\alpha]),f([\beta])\}$ also contained in $C(S)^{\iota}_{0}$. Since $f([\alpha])$ and $f([\beta])$ cannot have trivial geometric intersection, otherwise $f$ would not preserve the simplicial structure of ${\check{C}}(S)$, from Lemma 4.3, it follows that $f([\alpha])$ and $f([\beta])$ have geometric intersection $2$.
@@
We can then complete the two 00-simplices $\{[\alpha]\},\{[\beta]\}\in C(S)^{\iota}_{0}$ to two $(n(S)-4)$-simplices $v_{\alpha},v_{\beta}$ of $C(S)$ whose sets of vertices coincide except for the elements $[\alpha]\in v_{\alpha}$ and $[\beta]\in v_{\beta}$. In this way, we have defined an edge $\{v_{\alpha},v_{\beta}\}$ of the pants complex $C_{P}(S)\subset{\check{C}}_{P}(S)$ with the property that $\{f(v_{\alpha}),f(v_{\beta})\}$ is also an edge of ${\check{C}}_{P}(S)$. By Theorem 2.15, we conclude that $f\in\operatorname{Inn}({\widehat{\Gamma}}^{\pm}(S))$.
4.4. Proof of Theorem 1.1 for $S=S_{1,2}$
@@
By Lemma 2.16 and (ii) of Lemma 4.4, there is a natural monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5}))$ induced by restriction of automorphisms. Hence, by the case $S=S_{0,5}$ treated above, we have that $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))\subseteq\operatorname{Inn}({\widehat{\Gamma}}^{\pm}(S_{0,5}))$. This implies this case of the theorem, since, as in the proof of the case $S=S_{1,2}$ of Theorem 2.15, by Lemma 9.13 of [6], the inner automorphisms of ${\widehat{\Gamma}}^{\pm}(S_{0,5})$ which normalize its subgroup $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2})$ belong to $\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))$.
4.5. Proof of Theorem 1.1, for $g(S)\geq 1$ and $d(S)>2$
@@
Given an element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$, let us consider its action on ${\check{C}}(S)$. By Theorem 5.5 in [6], after possibly composing with an inner automorphism of $\mathrm{P}{\widehat{\Gamma}}(S)$, we may assume that $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$, where ${\hat{\mathrm{I}}}_{\gamma}$ is the procyclic subgroup associated to some nonseparating simple closed curve ${\gamma}$ on $S$. In particular, the automorphism $f$ preserves the star $\operatorname{Star}({\gamma})$ and, by Lemma 2.10, acts on the link $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{\gamma})$ through its image in the group $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{\gamma})$. Thus, $f$ preserves the vertex set of the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S)$, which identifies with profinite set of $(d(S)-1)$-simplices of $\operatorname{Star}({\gamma})$, and acts continuously on this set through an element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{\gamma})$. Since, by Lemma 2.8, $L_{\gamma}\cong{\check{C}}_{P}(S_{\gamma})$, the induction hypothesis implies that $f$ also preserves the edge set of $L_{\gamma}$. By Theorem 2.15, we then conclude that $f\in\operatorname{Inn}({\widehat{\Gamma}}^{\pm}(S))$.
@@
By Theorem 1.1, in order to prove Theorem 1.3, we have to show that, for $g(S)=0$, there holds $\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$. For this, we need to establish first that $\mathrm{P}{\widehat{\Gamma}}(S)$ is a characteristic subgroup of both $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and ${\widehat{\Gamma}}^{\pm}(S)$. We have:

### v3

@@ 4.1. Preliminary lemmas
For $g(S)=0$, the fixed-point set $C(S)_{0}^{\iota}$ of an antiholomorphic involution $\iota\in\mathrm{P}\Gamma(S)^{\pm}$ is finite and consists of isotopy classes of simple closed curves on $S$ which have between them geometric intersection either 0 or $2$.
Proof.
The complement $S\smallsetminus\operatorname{Fix}(\iota)$ is the disjoint union of two unpunctured discs. Let then $D^{\prime}$ and $D^{\prime\prime}$ be closed subdiscs of $S$ such that $D^{\prime}\cap D^{\prime\prime}=\operatorname{Fix}(\iota)$. Fix on the surface $S$ a hyperbolic complete metric compatible with the involution $\iota$. The geodesic representative $\alpha$ of $\{[\alpha]\}\in C(S)_{0}^{\iota}$ for this metric has then the property that $\alpha=\iota(\alpha)$ and the intersection $\alpha\cap D^{\prime}$ is a disjoint union of geodesic arcs with boundary in $\operatorname{Fix}(\iota)$. If $\alpha^{\prime}$ is one such arc, the union $\alpha^{\prime}\cup\iota(\alpha^{\prime})\subset\alpha$ is a simple closed curve, which immediately implies that $\alpha^{\prime}\cup\iota(\alpha^{\prime})=\alpha$. In particular, $\alpha$ crosses transversally $\operatorname{Fix}(\iota)$ in the two boundary points of $\alpha^{\prime}$. The isotopy class of $\alpha$ is then determined by the partition which $\alpha^{\prime}$ induces on the set of punctures of $S$ (which all lie on $\operatorname{Fix}(\iota)$). This implies the first statement of the proposition.
For $\{[\alpha]\}\neq\{[\beta]\}\in C(S)_{0}^{\iota}$, let also $\beta$ be the geodesic representative of the isotopy class $\beta]$ for the fixed metric. As above, we then have that $\beta=\iota(\beta)$ and the simple closed curve $\beta$ is the union of the two arcs $\beta^{\prime}=\beta\cap D^{\prime}$ and $\beta^{\prime\prime}=\beta\cap D^{\prime\prime}$. It is also clear that the geodesic arcs $\alpha^{\prime}$ and $\beta^{\prime}$ on the hyperbolic disc $D^{\prime}$ have geometric intersection either 0 or $1$. The second statement of the proposition then follows as well. ∎
By Lemma 4.3, there is a natural homomorphism $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$, for $d(S)>1$ and $S\neq S_{1,2}$, and a natural homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, for any hyperbolic surface $S$, both induced by restriction of automorphisms. We then have:
Lemma 4.7.
For $d(S)>1$, there holds:
@@
    For $S\neq S_{1,2}$, the homomorphism $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is injective.
@@
    The kernel of the homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ is trivial for $S\neq S_{2}$ and is generated by $\exp\zeta$ for $S=S_{2}$.
@@
(i): The statement is trivial for $n(S)\leq 1$ so that we can assume that either $g(S)=1$ and $n(S)>2$ or $g(S)\geq 2$ and $n(S)>1$. In particular, by Proposition 3.6, we can assume that the centralizer $Z_{{\check{\Gamma}}(S)^{\pm}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is trivial. This implies that the restriction of the homomorphism $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ to $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ is injective, so that the conclusion now follows from [9, Lemma 3.3].
(ii): Since $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))_{\mathrm{P}{\check{\Gamma}}(S)}$, in order to determine the kernel of the homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, we can use the Wells’ exact sequence associated to the short exact sequence $1\to\mathrm{P}{\check{\Gamma}}(S)\to\mathrm{P}{\check{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2\to 1$.
Since $\operatorname{Aut}(\mathbb{Z}/2)=\{1\}$, this implies that there is, in particular, a natural short exact sequence (cf. [9, Lemma 7.4]):
$1\to\operatorname{Hom}(\mathbb{Z}/2,Z(\mathrm{P}{\check{\Gamma}}(S)))\stackrel{{\scriptstyle\exp}}{{\to}}\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))_{\mathrm{P}{\check{\Gamma}}(S)}\to\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S)).$
If $S\neq S_{2}$, then $Z(\mathrm{P}{\check{\Gamma}}(S))=\{1\}$ and the conclusion follows. For $S=S_{2}$, we have $Z(\mathrm{P}{\check{\Gamma}}(S))\cong\mathbb{Z}/2$ and the only nontrivial element $\phi\in\operatorname{Hom}(\mathbb{Z}/2,Z(\mathrm{P}{\check{\Gamma}}(S)))$ is mapped to the automorphism $\exp(\phi)$ of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ defined by the assignment $x\mapsto x\cdot\phi(\mathrm{aug}(x))$, where $\mathrm{aug}\colon\thinspace\mathrm{P}{\check{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2$ is the orientation character. This coincides with the exceptional automorphism $\exp\zeta$ defined in Section 3.6. Hence, the second part of item (ii) follows. ∎
@@
For $d(S)>1$ and $S\neq S_{1,2}$, there is a chain of natural inclusions:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$
and, for $d(S)>1$, a natural homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, whose kernel is trivial for $S\neq S_{2}$ and is generated by $\exp\zeta$ for $S=S_{2}$.
@@
(i): This item immediately follows from Theorem 2.2, for the case $d(S)>1$, and from [9, Proposition 8.2] for the case $d(S)=1$.
(ii) and (iii): For $S=S_{1,2}$, by item (ii) of Proposition 3.11, there is a natural isomorphism $\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))\times\langle\exp\zeta\rangle$. For $S=S_{2}$, the subgroups $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ and $\langle\exp\zeta\rangle$ of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ are normal and have trivial intersection, so that they generate in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ a subgroup isomorphic to the direct product $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\times\langle\exp\zeta\rangle$.
@@
Therefore, by Lemma 4.8, in order to prove items (ii) and (iii) of Theorem 1.4, it is enough to show that the images of $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ coincide. The idea is to use Theorem 2.17. We need to consider separately three different cases.
4.3. Proof of Theorem 1.4, (ii), for $g(S)=0$
@@
By Lemma 4.5, after composing with an inner automorphism of ${\check{\Gamma}}(S)^{\pm}$, we can assume that a given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)^{\pm})$ preserves a fixed antiholomorphic involution $\iota\in\mathrm{P}\Gamma(S)^{\pm}$ whose fixed-point set in the surface $S$ is a separating simple closed curve containing all punctures of $S$.
@@
We will now make use of the results in Section 3.5. By Lemma 4.6, the fixed-point set $C(S)^{\iota}_{0}$, for the action of $\iota$ on $C(S)_{0}$, is finite and then, by Proposition 3.9, identifies with the fixed-point set ${\check{C}}(S)^{\iota}_{0}$, for the action of $\iota$ on ${\check{C}}(S)_{0}$.
@@
Let $\{\alpha,\beta\}$ be a pair of $\iota$-invariant simple closed curves on $S$ which intersect precisely in two points. An element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$, such that $f(\iota)=\iota$, then preserves $C(S)^{\iota}_{0}={\check{C}}(S)^{\iota}_{0}$ and sends the pair $\{[\alpha],[\beta]\}$ of 0-simplices in $C(S)^{\iota}_{0}$ to the pair $\{f([\alpha]),f([\beta])\}$ also contained in $C(S)^{\iota}_{0}$. Since $f([\alpha])$ and $f([\beta])$ cannot have trivial geometric intersection, otherwise $f$ would not preserve the simplicial structure of ${\check{C}}(S)$, from Lemma 4.6, it follows that $f([\alpha])$ and $f([\beta])$ have geometric intersection $2$.
@@
We can then complete the two 0-simplices $\{[\alpha]\},\{[\beta]\}\in C(S)^{\iota}_{0}$ to two (n(S)4)(n(S)-4)-simplices $v_{\alpha},v_{\beta}$ of $C(S)$ whose sets of vertices coincide except for the elements $\alpha]\in v_{\alpha}$ and $\beta]\in v_{\beta}$. In this way, we have defined an edge $\{v_{\alpha},v_{\beta}\}$ of the pants complex $C_{P}(S)\subset{\check{C}}_{P}(S)$ with the property that $\{f(v_{\alpha}),f(v_{\beta})\}$ is also an edge of ${\check{C}}_{P}(S)$. By Theorem 2.17, we conclude that $f\in\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$.
4.4. Proof of Theorem 1.4, (iii), for $S=S_{1,2}$
We will need an analogue, for extended mapping class groups, of Lemma 2.19:
Lemma 4.9.
The group $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5})$ identifies with an $\mathbb{I}$-characteristic subgroup of $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2})$. In particular, there is a natural homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5}))$ induced by restriction of automorphisms.
Proof.
As we remarked in Section 2.3, if we let $S_{/{\upsilon}}$ be the quotient of the surface $S_{1,2}$ by the hyperelliptic involution $\upsilon$} and let $B_{\upsilon}$ be the branch locus of the orbit map $S\to S_{/{\upsilon}}$, the surface $S_{/{\upsilon}}$ is a $1$-punctured sphere and there is a diffeomorphism $S_{/{\upsilon}}\smallsetminus B_{\upsilon}\cong S_{0,5}$. If we denote by $Q$ the puncture of $S_{0,5}$ which corresponds to the puncture of $S_{/{\upsilon}}$, there is then a natural isomorphism $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2})\cong{\check{\Gamma}}^{\pm}(S_{0,5})_{Q}$. In order to prove the lemma, it is then enough to prove that $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5})$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S_{0,5})_{Q}$.
Since $\mathrm{P}{\check{\Gamma}}(S_{0,5})$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S_{0,5})_{Q}$, the same argument of proof of Lemma 4.2 shows that there is a natural homomorphism $\Theta\colon\thinspace\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S_{0,5})_{Q})\to\Sigma_{B}$, where $B$ is a set of labels for the punctures of $S_{0,5}$.
The symmetric group $\Sigma_{B\smallsetminus Q}$ identifies with the stabilizer of $Q\in B$ for the tautological action of $\Sigma_{B}$ on $B$. The fact that $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{0,5})_{Q})$ is a normal subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S_{0,5})_{Q})$ and $\Theta(\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{0,5})_{Q}))=\Sigma_{B\smallsetminus Q}$ then implies that $\Theta(\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S_{0,5})_{Q}))=\Sigma_{B\smallsetminus Q}$ as well.
As in the proof of Lemma 4.3, we conclude that $\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5}))=\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{0,5})_{Q})\cap\ker\Theta$ is a normal subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S_{0,5})_{Q})$, which implies that $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5})$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S_{0,5})_{Q}$. ∎
By (ii) of Lemma 4.7 and Lemma 2.19, restriction of automorphisms induces a natural monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$. By Lemma 4.9, this homomorphism then factors through the natural homomorphisms $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5}))$ and $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5}))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$, also both induced by restriction of automorphisms.
@@
By the case $S=S_{0,5}$ of the theorem, treated above, we then have that the image of the group $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ is contained in the subgroup $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{0,5}))$. This implies this case of the theorem, since, as in the proof of Theorem 2.17 for $S=S_{1,2}$, by [9, Lemma 9.13], the inner automorphisms of ${\check{\Gamma}}^{\pm}(S_{0,5})$ which normalize the subgroup ${\check{\Gamma}}^{\pm}(S_{0,5})_{Q}\equiv\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2})$ belong to $\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))$.
4.5. Proof of Theorem 1.4, (ii) and (iii), for $g(S)\geq 1$ and $d(S)>2$
@@
Given an element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$, let us consider its action on ${\check{C}}(S)$. By [9, Theorem 5.5], after possibly composing with an inner automorphism of $\mathrm{P}{\check{\Gamma}}(S)$, we may assume that $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$, where ${\hat{\mathrm{I}}}_{\gamma}$ is the procyclic subgroup associated to some nonseparating simple closed curve $\gamma$} on $S$. In particular, the automorphism $f$ preserves the star $\operatorname{Star}({\gamma})$ and, by Lemma 2.12, acts on the link $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{\gamma})$ through its image in the group $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{\gamma})$. Thus, $f$ preserves the vertex set of the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S)$, which identifies with the profinite set of (d(S)1)(d(S)-1)-simplices of $\operatorname{Star}({\gamma})$, and acts continuously on this set through an element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{\gamma}))$. Since, by Lemma 2.9, $L_{\gamma}\cong{\check{C}}_{P}(S_{\gamma})$, the induction hypothesis implies that $f$ also preserves the edge set of $L_{\gamma}$. By Theorem 2.17, we then conclude that the image of $f$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ is contained in $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$.
5. Proof of Corollary 1.6
After possibly replacing ${\check{\Gamma}}(S)$ by ${\check{\Gamma}}(S)/Z({\check{\Gamma}}(S))$ (which is also isomorphic to a procongruence mapping class group), we can assume that the center $Z({\check{\Gamma}}(S))$ of ${\check{\Gamma}}(S)$ is trivial.
Let $f\in\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ be an automorphism which normalizes the subgroup $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$. This implies that $f$ satisfies the second condition of Lemma 2.18. Since the first one is satisfied by the above assumption and the third one is trivially satisfied, we have that $f$ extends to an automorphism of ${\check{\Gamma}}^{\pm}(S)$. The conclusion then follows from Theorem 1.4.
@@
By Theorem 1.4, in order to prove Theorem 1.7, we have to show that, for $g(S)=0$, there holds $\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$. For this, we need to establish first that $\mathrm{P}{\widehat{\Gamma}}(S)$ is a characteristic subgroup of both $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and ${\widehat{\Gamma}}^{\pm}(S)$. For this, we will need a series of lemmas.
The first part of the statement of the following lemma is well-known: